OpenStudy (anonymous):

A storage unit has the shape of a rectangular prism measuring 16 ft by 20 ft by 12 ft. What is the volume of this storage unit? A. 48 ft3 B. 288 ft3 C. 3840 ft3 D. 7680 ft3

OpenStudy (anonymous):

i pickC

OpenStudy (calculusxy):

\[V = l \times w \times h\] \[V = 16 \times 12 \times 20\]

OpenStudy (anonymous):

thx!

OpenStudy (calculusxy):

You're welcome :)